OpenStudy (anonymous):

Earthquakes release seismic waves that occur in concentric circles from the epicenter of the earthquake. Suppose a seismograph station determines the epicenter of an earthquake is located 9 kilometers from the station. If the epicenter is located at the origin, write the equation for the circular wave that passes through the station. A. x2 + y2 = 81 B. x2 + y2 = 9 C. (x - 9)2 + (y - 9)2 = 0 D. (x + 9)2 + (y + 9)2 = 0

OpenStudy (jamesj):

Which of these equations is the equation of circle of radius 9 centered at the origin?

OpenStudy (jamesj):

Hint: the equation of a circle centered at \( (a,b) \) and of radius \( r \) is \[ (x-a)^2 + (y-b)^2 = r^2 \]

OpenStudy (anonymous):

James <3 can you help me after by any chance?

OpenStudy (jamesj):

To make this as explicit as I can for you without giving the answer: if the circle is centered at the origin, then \[ (a,b) = (0,0) \] If the radius is 9 km, then \( r = 9 \). Now, using those values of the variables and the general form of the circle I wrote above, what is the equation of a circle, centered at the origin and of radius 9?

OpenStudy (jamesj):

\[ (x-a)^2 - (y-b)^2 = r^2 \] Now fill in the values of the variables, \(a, b\) and \( r \).

OpenStudy (jamesj):

**correction + \[ (x-a)^2 + (y-b)^2 = r^2 \]

OpenStudy (anonymous):

Okay so, Take the epicentre as the centre point of the circle (Since earthquake waves are circular.) Then, since the epicentre is 9 km away, let the station be any point on the circle. So, radius= 9^2 = 81 So, equation of the circle: x^2 + y^2 = 81 is your answer.